Circular No. 6145

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995H IN NGC 3526 J. Mueller reports her discovery of a supernova of mag ∼ 16 in NGC

3526 (α = 11^{h}07^{m}.0, δ = +7^{o}10′, equinox 2000.0), found on an IVN plate taken on Feb. 24 in the course of the second Palomar Sky Survey by Mueller, D. Griffith, and J. D. Mendenhall; SN 1995H is 17′′ west and 3′′ south of the galaxy’s center. S. Benetti, European Southern Observatory (ESO); M. Longhetti, Osservatorio Astronomico di Brera; and G. Marconi, Osservatorio Astronomico, Rome, report that preliminary inspection of a fullyreduced CCD spectrogram (range 400–900 nm; resolution 0.6 nm) obtained on Mar. 7.17 UT with ESO 1.5-m telescope (+ Boller & Chivens spectrograph), shows that the object located 19′′.6 west and 7′′.5 south of the galaxy’s center is a type II supernova, 6–7 weeks after outburst. Strong P-Cyg lines of H, Fe II (multiplets 40 and 42), Sc II, Na I (D), O I (777.3 nm), and Ca II (infrared triplet) are superimposed on a relatively red continuum (∼ 5800 K). The expansion velocity deduced from the minima of the Hβ and Hα lines is ∼ 7300 km/s. CCD magnitude obtained of SN 1995H on Mar. 7.21 with 1.54-m Danish telescope at La Silla: V = 16.0. P. Garnavich, Center for Astrophysics, confirms the type-II classification from a spectrogram obtained by J. Peters with the Mt. Hopkins 1.5-m Tillinghast telescope on March 7.42; he notes that there is significant contamination from H II emission in the host galaxy. R. Stathakis, Anglo-Australian Observatory; and B. Schmidt, Mount Stromlo and Siding Spring Observatories, report that their spectrogram (range 320–900 nm) obtained on Mar. 7.56 at the Anglo-Australian Telescope suggests that the type-II supernova is ∼ 3 weeks after explosion. They also note the P-Cyg profiles of Hα, Ca II, and Fe II; the expansion rate of the photosphere, derived from Fe II (516.9 and 501.8 nm), is 7000 km/s, and the redshift of the host galaxy, measured from narrow Hα at the position of SN 1995H, is 1450 km/s.

NOVA IN THE LARGE MAGELLANIC CLOUD 1995 Johnson UBV and Cousins RI photometry: Mar. 3.496 UT, V = 11.26,

U − B = −0.42, B − V = +0.30, V − R = +0.22, V − I = +0.47 (A. C. Gilmore, Mt. John University Observatory; through cirrus; comparison stars Cousins CPD −68^{o}311 and −65^{o}443); 4.417, 11.42, −0.57, +0.20, +0.27, +0.54 (Gilmore); 5.393, 11.75, −0.61, +0.28, –, – (G. W. Christie, Auckland Observatory; comparison stars HD 37027 and HD 36584); 5.401, 11.75, −0.57, +0.23, –, – (Christie).

1995 March 7 Daniel W. E. Green